# Rossby Wave Instability and Substructure Formation in 3D Non-Ideal MHD Wind-Launching Disks

### v1

@@ Abstract
The RWI occurs when the local Rossby wave is trapped in the non-self-gravitating disks. It may allow non-axisymmetric perturbations to grow into well-defined vortices (e.g., Zaqarashvili et al., 2021). The RWI has been investigated in the 2D (vertically integrated) linear theories (e.g., Lovelace et al., 1999; Li et al., 2000; Umurhan, 2010; Ono et al., 2016) and numerical simulations (e.g., Li et al., 2001; Varnière & Tagger, 2006; Lyra et al., 2008; Ono et al., 2018; Cimerman & Rafikov, 2023). In the 3D disk (where the vertical structure is accounted for), the RWI has also been studied using both linear theories (e.g., Lin, 2012, 2013) and numerical simulations (e.g., Lyra & Mac Low, 2012; Richard et al., 2013), mostly with a prescribed initially axisymmetric substructure that is observationally motivated but not formed self-consistently. The spontaneous formation of rings and gaps in 2D (axisymmetric) non-ideal MHD simulations (e.g., Suriano et al., 2017, 2018; Nolan et al., 2023) provides a rare opportunity to examine the RWI of axisymmetric structures produced through consistent dynamics using 3D simulations. Just as importantly, such 3D simulations also allow us to investigate the formation of disk substructures and the development of RWI in them simultaneously. In this paper, we seek to determine whether the non-linear development of RWI in the rings and gaps formed in the non-ideal MHD wind-launching disks can lead to the formation of vortices and, if so, how they affect the rings and gaps in the radial gas distribution. Vortices are important to investigate because of their potential for dust trapping (see, e.g., a recent review by Bae et al., 2023), which may facilitate planetesimal formation.
@@ 4.3 RWI Vortex Structure
An important parameter to characterize a vortex is its aspect ratio , where and are the major and minor axes of close streamlines (e.g., Ono et al., 2018). Because the velocity field in our non-ideal MHD simulation is much more disordered than the hydro simulations (see the velocity streamline in Fig. 7), we choose to approximate the shape of the vortex using the iso-density contour of (i.e., the black lines for the dominant vortex in Fig. 7a). The resulting aspect ratio for the dominant vortex near the 8 au radius is shown in Fig. 9 as a function of time during the period when the mode dominates (; see panel b of Fig. 5). The aspect ratio is of order 10 or larger for the entire duration, indicating that the vortex is highly elongated, consistent with visual inspection of Fig. 7 and especially Fig. 4 and their associated animations. Lesur & Papaloizou (2009) pointed out that vortices in disks tend to be unstable to ellipsoidal instability in 3D, although Richard et al. (2013) found that highly elongated vortices with or larger can survive in 3D compressible hydro simulations. This hydro result contrasts with our non-ideal MHD simulation, where the amplitude of the dominant mode at 8 au steadily decreases after peaking around despite its large aspect ratio.
@@ 5 Conclusion
An interesting question to address in future investigations is how RWI affects the dust distribution and the disk appearance in continuum emission. Most observed disk dust substructures appear rather axisymmetric, although a small fraction shows non-axisymmetric features, including arcs, which are potentially vortices. The RWI vortices discussed in this paper open up the possibility of producing such dusty vortices in non-ideal MHD wind-launching disks. However, detailed dust treatment is needed to explore this possibility quantitatively. The current simulation will also benefit from adaptive (rather than static) mesh refinement, which will better resolve the vortices formed in the global 3D simulations.
@@ Acknowledgements
CYH acknowledges support from NASA SOFIA grant SOF-10_0505 and NRAO ALMA Student Observing Support (SOS) and computing resources from UVA research computing (RIVANNA) and NASA High-Performance Computing. ZYL is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. MKL is supported by the National Science and Technology Council (grants 111-2112- M-001-062-, 112-2112-M-001-064-, 111-2124-M-002-013- , 112-2124-M-002-003-) and an Academia Sinica Career Development Award (AS-CDA110-M06).
@@
The inclusion of a Data Availability Statement is a requirement for articles published in MNRAS. Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article. The statement may refer to original data generated in the course of the study or to third-party data analysed in the article. The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs.

### v2

@@ Abstract
The RWI occurs when the local Rossby wave is trapped in the non-self-gravitating disks. It may allow non-axisymmetric perturbations to grow into well-defined vortices (e.g., Zaqarashvili et al., 2021). The RWI has been investigated in the 2D (vertically integrated) linear theories (e.g., Lovelace et al., 1999; Li et al., 2000; Umurhan, 2010; Ono et al., 2016) and numerical simulations (e.g., Li et al., 2001; Varnière & Tagger, 2006; Lyra et al., 2008; Ono et al., 2018; Cimerman & Rafikov, 2023). In the 3D disk (where the vertical structure is accounted for), the RWI has also been studied using both linear theories (e.g., Lin, 2012, 2013) and numerical simulations (e.g., Lyra & Mac Low, 2012; Richard et al., 2013), mostly with a prescribed initially axisymmetric substructure that is observationally motivated but not formed self-consistently. RWI has also been studied in magnetized disks, with toroidal magnetic fields (e.g., Yu & Li, 2009), large-scale poloidal magnetic fields (e.g., Yu & Lai, 2013), and non-ideal MHD effects (e.g., Cui et al., 2024). The spontaneous formation of rings and gaps in 2D (axisymmetric) non-ideal MHD simulations (e.g., Suriano et al., 2017, 2018; Nolan et al., 2023) provides a rare opportunity to examine the RWI of axisymmetric structures produced through consistent dynamics using 3D simulations. Just as importantly, such 3D simulations also allow us to investigate the formation of disk substructures and the development of RWI in them simultaneously. In this paper, we seek to determine whether the non-linear development of RWI in the rings and gaps formed in the non-ideal MHD wind-launching disks can lead to the formation of vortices and, if so, how they affect the rings and gaps in the radial gas distribution. Vortices are important to investigate because of their potential for dust trapping (see, e.g., a recent review by Bae et al., 2023), which may facilitate planetesimal formation.
@@ 4.3 RWI Vortex Structure
An important parameter to characterize a vortex is its aspect ratio , where and are the major and minor axes of close streamlines (e.g., Ono et al., 2018). Because the velocity field in our non-ideal MHD simulation is much more disordered than the hydro simulations (see the velocity streamline in Fig. 7), we choose to approximate the shape of the vortex using the iso-density contour of (i.e., the black lines for the dominant vortex in Fig. 7a). The resulting aspect ratio for the dominant vortex near the 8 au radius is shown in Fig. 9 as a function of time during the period when the mode dominates (; see panel b of Fig. 5). The aspect ratio is of order 10 or larger for the entire duration, indicating that the vortex is highly elongated, consistent with visual inspection of Fig. 7 and especially Fig. 4 and their associated animations. Lesur & Papaloizou (2009) pointed out that vortices in disks tend to be unstable to ellipsoidal instability in 3D, although the maximum growth rate obtained from linear analysis decreases rapidly with increasing aspect ratio in a stratified disk (see the red curve in their Fig. 6), making it increasingly difficult to detect in simulations. Richard et al. (2013) found that highly elongated vortices with or larger can survive in 3D compressible hydro simulations. This hydro result contrasts with our non-ideal MHD simulation, where the amplitude of the dominant mode at 8 au steadily decreases after peaking around despite its large aspect ratio.
@@
A caveat of our simulations is that, despite the use of static mesh refinement, the resolution may not be high enough to fully resolve any MRI turbulence that may develop, especially in the relatively well (magnetically) coupled outer part of the disk beyond au, where the initial ambipolar Elsasser number on the midplane exceeds 10 for the 3D models discussed in this section (see the red curve in the upper panel of Fig. 16). Although our focus is on the less well-coupled inner disk (inside au) that is expected to be less prone to MRI than the outer disk, it is possible that its dynamics and structure can be modified at a higher resolution, where a better resolved MRI turbulence can potentially reduce the amplitudes of the rings and gaps and increase their widths, making them more stable to RWI. Indeed, there are no obvious vortices in the 3D model shown in Fig. 10 of Cui and Bai (2021), which has a higher resolution than ours but a smaller azimuthal angle range ( rather than ). Whether the limited azimuthal angle range has a strong effect on the formation and evolution of RWI vortices or not remains to be determined. Higher-resolution simulations with a full azimuthal angle range are desirable to resolve the potential discrepancy.
@@ 5 Conclusion
An interesting question to address in future investigations is how RWI affects the dust distribution and the disk appearance in continuum emission. Most observed disk dust substructures appear rather axisymmetric, although a small fraction shows non-axisymmetric features, including arcs, which are potentially vortices. The RWI vortices discussed in this paper open up the possibility of producing such dusty vortices in non-ideal MHD wind-launching disks. However, detailed dust treatment is needed to explore this possibility quantitatively. The current simulation will also benefit from adaptive (rather than static) mesh refinement, which will better resolve the vortices formed in the global 3D simulations. Higher resolution is also desirable to better resolve any MRI turbulence that may develop, which can modify the disk dynamics and structure, which, in turn, affects its stability to RWI.
@@ Acknowledgements
We thank the referee, Dr. Cui, Can, for a prompt and constructive report. CYH acknowledges support from NASA SOFIA grant SOF-10_0505 and NRAO ALMA Student Observing Support (SOS) and computing resources from UVA research computing (RIVANNA) and NASA High-Performance Computing. ZYL is supported in part by NASA 80NSSC20K0533 and NSF AST-2307199. MKL is supported by the National Science and Technology Council (grants 111-2112- M-001-062-, 112-2112-M-001-064-, 111-2124-M-002-013- , 112-2124-M-002-003-) and an Academia Sinica Career Development Award (AS-CDA110-M06).
@@
The data underlying this article will be shared on reasonable request to the corresponding author.